OpenStudy (anonymous):

Just want to make sure i am doing this right. Write an equation in point-slope form for the line through the given point with the given slope. (8,3): m=6 I got the answer y-3=6(x-8) is that correct?

OpenStudy (anonymous):

correct u can simplify further

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

Bring the variables to one side

OpenStudy (anonymous):

how do i bring the varibles to one side?

OpenStudy (anonymous):

simplify both sides and then arrange

OpenStudy (anonymous):

Since you have done most of the work the final answer is y=6x-45

OpenStudy (anonymous):

y-3=6x -48 or 0=6x -48 -y +3 or 6x-y -45 =0 or 6x+y=45

OpenStudy (anonymous):

@matricked 6x+y=45 is wrong. you maade a mistake while rearranging

OpenStudy (anonymous):

@Varunkumar6 thanks it is 6x-y=45